EXCESS CAPACITY

Krogh Lumber's 2016 financial statements are shown here.

Krogh Lumber: Balance Sheet as of December 31, 2016 (Thousands of Dollars)
Cash $1,800 Accounts payable $7,200
Receivables 10,800 Notes payable 3,472
Inventories 12,600 Accrued liabilities 2,520
Total current assets $25,200 Total current liabilities $13,192
Mortgage bonds 5,000
Net fixed assets 21,600 Common stock 2,000
Retained earnings 26,608
Total assets $46,800 Total liabilities and equity $46,800


Krogh Lumber: Income Statement for December 31, 2016 (Thousands of Dollars)
Sales $36,000
Operating costs including depreciation 30,783
Earnings before interest and taxes $5,217
Interest 1,017
Earnings before taxes $4,200
Taxes (40%) 1,680
Net income $2,520
Dividends (60%) $1,512
Addition to retained earnings $1,008
  1. Assume that the company was operating at full capacity in 2016 with regard to all items except fixed assets; fixed assets in 2016 were being utilized to only 61% of capacity. By what percentage could 2017 sales increase over 2016 sales without the need for an increase in fixed assets? Round your answer to two decimal places.
    %

  2. Now suppose 2017 sales increase by 30% over 2016 sales. Assume that Krogh cannot sell any fixed assets. All assets other than fixed assets will grow at the same rate as sales; however, after reviewing industry averages, the firm would like to reduce its operating costs/sales ratio to 80% and increase its total liabilities-to-assets ratio to 42%. The firm will maintain its 60% dividend payout ratio, and it currently has 1 million shares outstanding. The firm plans to raise 35% of its 2017 forecasted interest-bearing debt as notes payable, and it will issue bonds for the remainder. The firm forecasts that its before-tax cost of debt (which includes both short- and long-term debt) is 10.5%. Any stock issuances or repurchases will be made at the firm's current stock price of $40. Develop Krogh's projected financial statements. What are the balances of notes payable, bonds, common stock, and retained earnings? Round your answers to the nearest hundredth of thousand of dollars.
    Krogh Lumber Pro Forma Income Statement December 31, 2017 (Thousands of Dollars)
    2016 2017
    Sales $36,000 $
    Operating costs (includes depreciation) 30,783
    EBIT $5,217 $
    Interest expense 1,017
    EBT $4,200 $
    Taxes (40%) 1,680
    Net Income $2,520 $
    Dividends $1,512 $
    Addition to RE $1,008 $
    Krogh Lumber Pro Forma Balance Statement December 31, 2017 (Thousands of Dollars)
    2016 2017
    Assets
    Cash $1,800 $
    Accounts receivable 10,800
    Inventories 12,600
    Fixed assets 21,600
    Total assets $46,800 $
    Liabilities and Equity
    Payables + accruals $9,720 $
    Short-term bank loans 3,472
      Total current liabilities $13,192 $
    Long-term bonds 5,000
      Total liabilities $18,192 $
    Common stock 2,000
    Retained earnings 26,608
      Total common equity $28,608 $
    Total liab. and equity $46,800 $
Answer #1
1..Full capacity sales=36000/61%=
59016.39
So,
Percentage by which 2017 sales could increase over 2016 sales without the need for an increase in fixed assets=
(59016.39-36000)/36000=
63.93%
c.. Krogh Lumber Pro Forma Income Statement December 31, 2017 (Thousands of Dollars)
2016 2017
Sales $36,000 36000*1.3= 46800
Operating costs (includes depreciation) 30,783 46800*80%= 37440
EBIT $5,217 9360
Interest expense 1,017 (3568.32+6626.88)*10.5% 1070.50
EBT $4,200 8289.50
Taxes (40%) 1,680 8289.5*40%= 3315.80
Net Income $2,520 4973.70
Dividends $1,512 4973.70*60%= 2984.22
Addition to RE $1,008 1989.48
Krogh Lumber Pro Forma Balance Statement December 31, 2017 (Thousands of Dollars)
2016 2017
Assets
Cash $1,800 1800*1.3= 2340
Accounts receivable 10,800 10800*1.3= 14040
Inventories 12,600 12600*1.3= 16380
Fixed assets 21,600 21600
Total assets $46,800 54360
Liabilities and Equity
Payables + accruals $9,720 9720*1.3= 12636
Short-term bank loans 3,472 (22831.2-12636)*35%= 3568.32
  Total current liabilities $13,192 16204.32
Long-term bonds 5,000 (22831.2-12636)*65%= 6626.88
  Total liabilities $18,192 54360*42%= 22831.20
Common stock 2,000 Tot. common eq.-Ret. Earn. 2931.32
Retained earnings 26,608 26608+1989.48= 28597.48
  Total common equity $28,608 Total Liab.& equity-Tot.liab.--(54360-22831.20) 31528.80
Total liab. and equity $46,800 54360
ANSWER:
Balances of :  
(Rounded to the nearest hundredth of thousand of dollars)
Notes payable 3600
Bonds 6600
Common stock 2900
Retained earnings 28600